a canoeist paddled for 4 hours with a 6km/h current and travels a distance of 80 km. how fast can the canoeist paddle in still water?
Answer by mananth(16949)

d= 60 km
let paddling speed in still water be x km/h
with current speed = (x+6) km/h
t= 4 hours
D= r*t
160=(x+6)*4
divide by 4
40= x+6
subtract 6 from both sides
34=x
34 km/h in still water